# Earth Observation Satellite Scheduling with Graph Neural Networks

### v1

@@ 4 Solution
Figure 2: General Architecture
Following [11], we use a reinforcement learning setup where the agent receives continuous graphs representing partial schedules as input, selects the next observations to schedule, and updates its parameters based on the reward representing the cumulative utility of the schedules it produces. For a given problem, a simulator is in charge of managing the different graphs and feeding the learner with the appropriate data. The learner implements a policy, that is, a stochastic mapping from states $s$ to actions $a$ as defined above. It learns a policy that maximizes the reward function over a base of real-world problems used as training set. The policy has to be able to generalize to test problems, that is, exhibit good performances without further learning on a set of instances not seen before.
@@ 4.1 Reinforcement Learning
As our core algorithm, we use the Proximal Policy Optimization (PPO) algorithm [16] with action masking [10].
A peculiar aspect of the EOSP instances we have to solve is that the utility of different acquisitions may vary by up to 8 degree of magnitude. In fact, acquisitions are grouped in 7 priority classes with utility value ranging from 1 to $10^{8}$. The utility of the observations within a class of priority is equal to the value of that class, plus a small term depending on the predicted cloud coverage at the location of the acquisition (in order to favor acquisitions that are likely to happen with a clear sky). This generates instability in DRL algorithms (and in MDPs in general), as the low priority observations provide a reward that might be difficult to distinguish form the noise in the algorithm. In addition, the critic must learn very large values, starting from very low values at initialization, and following tiny gradient steps. This makes learning slow and inefficient.
We tried several approaches to handle this, including using a logarithmic scale and 2-hot encoding [7]. In our current implementation, we simply divide each individual reward by the average utility of all candidates observations in the problem. This is a simple way to remedy the issue of having to learn very large values, but it does not fix the problem of the discrepancy between rewards (unless some extreme priority classes are not represented in the problem instance). We are currently examining optimization with lexicographic preferences [17].
@@ 4.2 GNN Implementation
Node attributes:
To inform the learner, we label each node of the continuous graph with the window of visibility of the corresponding observation. We note that the continuous graph still does not contain any information about the transitions duration, so the learner would be blind to them. To compensate for this, each node $i$ of the continuous graph is labeled with information about the satellite attitude while performing observation $i$, namely, the min, max and average pitch and roll of the satellite over the observation VTW. Although this information is not sufficient to recover the exact duration of transitions, it allows the learner inferring them closely enough to perform well, as shown in our simulation results.
Graph rewiring:
A Message-Passing Graph Neural Network (MP-GNN) [23] uses a graph structure as a computational lattice. It propagates information, represented as messages, along the oriented graph edges only. In our case, if an MP-GNN uses only the EOSP continuous graph edges, then we explicitly forbid information to flow from future acquisitions to the present choice of the next acquisition. This is definitely not what we want: we want the agent to choose the next observation to schedule based on its effect on the on future conflicts. In other words, we want information to go from future to present tasks. Therefore, we have to edit the input graph before it can used by the MP-GNN. This is known in the MP-GNN literature as “graph rewiring”.
For every (precedence) edge in the continuous graph, a link pointing in the other direction is added to the rewired graph (reverse-precedence). Different edge types are defined for precedence and reverse-precedence edges, to enable the learned operators to differentiate between chronological and reverse-chronological links. The systems learns to pass information in a forward and backward way, depending on what is found useful during learning.
Embeddings:
@@
Graph pooling:
A node is added and connected to every other node to allow collecting global information about the entire graph, as opposed to the local information associated with the nodes of the original graph. It is used by the critic to estimate the value of the graph as a whole. It is also used by the actor, where the global graph encoding is concatenated to each node embedding. Indeed, messages are passed by the MP-GNN algorithm only between immediate neighbors. Therefore, a network of depth n_layersn\_layers is able to anticipate only n_layersn\_layers observations ahead. Having the global node embedding concatenated to each node embedding compensates for this, allowing the current decision to take into account the entire graph. [^1] Adding such kind of node to the graph is equivalent to learning a custom graph pooling operator.
GNN:
As a message passing GNN, we use EGATConv from the DGL library [21], which enriches GATv2 graph convolutions [2] with edges attributes. We used 4 attention heads, leading to an output of size $4\timeshidden$_dim. This dimension is reduced to hidden_dim using learnable MLPs, before being passed to next layer (in the spirit of feed-forward networks used in transformers [20]). The output of a layer can be summed with its input using residual connections [8]. For most of our experiments, we used 10 such layers. The message-passing GNN yields a value for every node, and a global value for the graph (from the graph pooling node).
Action selection:
@@
Dealing with different problem sizes:
@@
Connecting to PPO:
In most generic PPO implementation, the actor (policy) consists of a feature extractor whose structure depends on the data type of the observation, followed by a MLP whose output dimension matches the number of actions. The same holds for the critic (value estimator), with the difference that the output dimension is 1. Some layers can be shared (the feature extractor and first layers of the MLPs). In our case, we do not want to use such a generic structure because we have a one-to-one matching from the number of nodes to the actions. We thus always keep the number of nodes as a dimension of the data tensors.
@@ 5 Experiments
We use a set of real-life problems provided by an industrial partner who owns Earth observation satellites. As explained in Section 3, problems are given in the form of a discrete-time graph. The simulator uses this graph to compute and maintain the continuous-time graph. To provide intuition on the difficulty of the problem, Table 1 shows some statistics on a few test problems and their representation as graphs.
@@
# Acquisitions Discrete graph Continuous graph Ratio Discrete Continuous Ratio
@@
Table 1: Exemple problem sizes
We compare our DRL approach to two solutions currently being used for operating such satellites.
Greedy algorithm:
It is the algorithm that is currently used for real-world operations. It greedily selects acquisitions to add to the schedule based on the utility, and inserts them in the plan if possible. Previously selected tasks may be slightly postponed, but never canceled.
RAMP:
[1] It is an implementation of a Dijkstra search algorithm in the discrete graph. Although based on an admissible algorithm (Dijkstra), RAMP is not guaranteed to find the absolute optimal schedule. This is due to the exclusion links between nodes of the discrete graph that represents the same task started at different times. Nevertheless, RAMP constantly provides the best known schedules on real problems. Unfortunately, its complexity prevents using it for real-time operations. Therefore, it is used as a reference in these simulation results.
@@ 5.1 Unitary Score
First, we compare our approach to baselines on a relaxed problem: we try to maximize the number of acquisitions scheduled, irrespectively of their priority or utility. This measure of performance is insensitive to the large gaps in acquisitions utility discussed in Section  4.1. We run two experiments:
Single problem:
First, we want to measure if our models and algorithms can possibly achieve competitive performance on a given problem. We train our learner on a single problem with a total of 106 acquisitions and let it overfit as much as it needs, as long as it achieves great performance. As illustrated in Fig. 3-left, we observe that it is indeed able to beat both the greedy algorithm and RAMP scores. This result shows that our architecture is able to implement very powerful policies. In the next set of experiments, we put it to the challenge of realistic learning environment.
Figure 3: Unitary scores. Left: Training on a single problem of 106 acquisitions. Right: Test performance on 31 unseen problems, after training on 128 different problems.
Generalization:
To measure the ability to transfer knowledge from one task to another, we train on 128 problems of about 100 acquisitions and test on 31 unseen problems of similar size. The learning curve of Fig. 3-right shows the evolution of the performance on the test set, as learning progresses. We also measure the number of times where Wheatley’s performances are above, below or equal to the greedy algorithm and RAMP (Fig. 4) on the test problems. This shows that our system is able to generalize to unseen problems, outperforming the currently deployed solution.
Figure 4: Unitary scores for Wheatley vs. Greedy (left) and vs; Ramp (right) on 31 test problems.
@@ 5.2 General Utility
In our second sets of experiments, we take into account the utility of observations and aim at maximizing the cumulative utility of all the observations included in the final schedule, as in the full-fledged MDP framework presented in Section 7. As before, we perform two sets of experiments: one where the learner is free to overfit on a single problem to reach its best performances, and one aiming at measuring its ability to generalize.
Single Problem:
Our test on single problem with a total of 88 candidate acquisitions shows that our system is able to outperform the greedy algorithm and reach the score of RAMP ( Figure 5-left). This proves the suitability of the architecture for the full MDP set up.
Figure 5: Utilities obtained when training on a single problem with 88 acquisitions (left) and averaged on 27 unseen problems, after a training on 639 different problems (right).
Generalization:
We train on 639 problems of about 100 acquisitions and test on 27 unseen problems of similar size. The learning curves are displayed in Figure 5-right and show that the learner is able to generalize. The plot showing the number of times where Wheatley is above, below or equal to the greedy solution and RAMP are presented in Fig. 6. We see that Wheatley outperforms the deployed solution and approaches the best known performances, in a realistic set-up where problems are not known in advance. Fig. 7 shows examples of solutions produced by Greedy and Wheatley. We can see that Wheatley finds a smoother and more efficient trajectory for the satellite.
Table 2 shows comprehensive results for the agent trained on problems of size 100, evaluated on different sizes of problems with operational scores. The last line is an evaluation on instances with many conflicts, where RAMP performs worse than the greedy algorithm. Results show that Wheatley performs very well on not too large instances but is quite outperformed by the greedy approach on the largest instance. However, it is quite competitive in the case of highly conflictual instances, which is promising for future works.
Figure 6: Utilities for Wheatley vs. Greedy (left) and vs. Ramp (right) on 27 test problems.
@@ 6 Conclusion
We showed that DL-based approaches to the EOSP are challenging some of the best known techniques. There are several perspectives we are currently exploring to extend this work. First, as stated before, we are trying to take advantage of the large discrepancy in acquisition utility by using lexicographic RL algorithms such as [17]. Scheduling tasks by decreasing priority would provide stronger guarantees to find the optimal schedule. To achieve this, schedules must be built in a non-chronological order, which is not the case in our current implementation. Currently, we choose the next acquisition to insert just after the last inserted one, using some foresight given by the GNN. This foresight is limited by the number of layers of the GNN. As we said, the discrete-time graph is tailored for state-space search and chronological insertion. Future work will consider developing an alternative continuous-tine graph representation of the EOSP where observations can be added to the schedule in any order, using Simple Temporal Networks [5]. Such work will open promising avenues for using lexicographic preferences.

### v2

@@ 4 Solution
Figure 2: Wheatley general architecture
Following (wheatley-jssp), we use a reinforcement learning setup where the agent receives continuous graphs representing partial schedules as input, selects the next observations to schedule, and updates its parameters based on the reward representing the cumulative utility of the schedules it produces. For a given problem, a simulator is in charge of managing the different graphs and feeding the learner with the appropriate data. The learner implements a policy, that is, a stochastic mapping from states $s$ to actions $a$ as defined above. It learns a policy that maximizes the reward function over a base of real-world problems used as training set. The policy has to be able to generalize to test problems, that is, exhibit good performances without further learning on a set of instances not seen before.
@@ 4.1 Reinforcement Learning
As our core algorithm, we use the Proximal Policy Optimization (PPO) algorithm (PPO) with action masking (ppo_mask), due to its relative simplicity and its good results on many different problems.
A peculiar aspect of the EOSP instances we have to solve is that the utility of different acquisitions may vary by up to 8 degrees of magnitude. In fact, acquisitions are grouped in 7 priority classes with utility value ranging from 1 to $10^{8}$. The utility of the observations within a class of priority is equal to the value of that class, plus a small term depending on the predicted cloud coverage at the location of the acquisition (in order to favor acquisitions that are likely to happen with a clear sky). This generates instability in DRL algorithms (and in MDPs in general), as the low priority observations provide a reward that might be difficult to distinguish from noise in the algorithm. In addition, the critic must learn very large values, starting from very low values at initialization, and following tiny gradient steps. This makes learning slow and inefficient.
We tried several approaches to handle this, including using a logarithmic scale and 2-hot encoding (dreamerv3). In our current implementation, we simply divide each individual reward by the average utility of all candidate observations in the problem. This is a simple way to remedy the issue of having to learn very large values, but it does not fix the problem of the discrepancy between rewards (unless some extreme priority classes are not represented in the problem instance). We are currently examining optimization with lexicographic preferences (lexicoRL).
@@ 4.2 GNN Implementation
Node attributes
To inform the learner, we label each node $i$ of the continuous graph $\mathcal{G}^{C}$ with the visibility window $W_{i}$. The continuous graph does not bring any information about the transition duration to the learner. To compensate for this, each node $i$ of the continuous graph is labeled with information about the satellite attitude while performing observation $i$, namely, the min, max and average pitch and roll angles of the satellite over the observation VTW. Although this information is not sufficient to recover the exact duration of transitions, it allows the learner to infer them closely enough to perform well, as shown in our simulation results.
Graph rewiring
A Message-Passing Graph Neural Network (MP-GNN) (gin) uses a graph structure as a computational lattice. It propagates information, represented as messages, along the oriented graph edges only. In our case, if an MP-GNN uses only the EOSP continuous graph edges, then information cannot flow from future acquisitions to the present choice of the next acquisition. This is definitely not what is desired: the agent should choose the next observation to schedule based on its effect on future conflicts. In other words, we want information to go from future to present tasks. Therefore, we edit the input graph before it can be used by the MP-GNN. This is known in the MP-GNN literature as “graph rewiring”.
For every (precedence) edge in the continuous graph, a link pointing in the other direction is added to the rewired graph (reverse-precedence). Different edge types are defined for precedence and reverse-precedence edges, to enable the learned operators to differentiate between chronological and reverse-chronological links. The system learns to pass information in a forward and backward way, depending on what is found useful during learning.
Embeddings
@@
Graph pooling
A node is added and connected to every other node to allow collecting global information about the entire graph, as opposed to the local information associated with the nodes of the original graph. It is used by the critic to estimate the value of the graph as a whole. It is also used by the actor, where the global graph encoding is concatenated to each node embedding. Indeed, messages are passed by the MP-GNN algorithm only between immediate neighbors. Therefore, a network of depth n_layersn\_layers is able to anticipate only n_layersn\_layers observations ahead. Having the global node embedding concatenated to each node embedding compensates for this, allowing the current decision to take into account the entire graph. [^3] Adding such a kind of node to the graph is equivalent to learning a custom graph pooling operator.
GNN
As a message passing GNN, we use EGATConv from the DGL library (dgl), which enriches GATv2 graph convolutions (gatv2) with edge attributes. We used 4 attention heads, leading to an output of size $4\timeshidden$_dim. This dimension is reduced to hidden_dim using learnable MLPs, before being passed to the next layer (in the spirit of feed-forward networks used in transformers (attention)). The output of a layer can be summed with its input using residual connections (residual). For most of our experiments, we used 10 such layers. The message-passing GNN yields a value for every node and a global value for the graph (from the graph pooling node).
Action selection
@@
Dealing with different problem sizes
@@
Connecting to PPO
In most generic PPO implementations, the actor (policy) consists of a feature extractor whose structure depends on the data type of the observation, followed by a MLP whose output dimension matches the number of actions. The same holds for the critic (value estimator), with the difference that the output dimension is 1. Some layers can be shared (the feature extractor and first layers of the MLPs). In our case, we do not want to use such a generic structure because we have a one-to-one matching from the number of nodes to the actions. We thus always keep the number of nodes as a dimension of the data tensors.
4.3 Inference Time Search
Ideally, the agent should be able to select the best action simply by selecting the argmax of the scores of the candidate actions. But several reasons may lead to suboptimal behavior doing so: the learning may not be finished (for instance if a plateau is reached during learning phase), the GNN may not be able to retain information far enough into the future of the current schedule (due to GNN over-smoothing or over-squashing, and due to the fact that we choose actions chronologically), or the agent may not have enough generalization ability (it learns on a given set of scenarios, and it is difficult to measure the closeness of these setups to evaluation/real-world setups). For all these reasons, we propose to perform search at inference time, meaning after the learning phase of the agent, in order to further refine the policy. Several well-known techniques are possible, like beam search or Monte-Carlo Tree Search. We found beam search to be hard to evaluate because of the very large branching factor we face, leading to exploring very few beams in reasonable time/memory. We thus conducted some experiments with MCTS, using PUCT bound (Silver_2016).
For doing so, we allocate some budget for MCTS exploration. One key aspect of MCTS is the way to set the initial value of expanded nodes; in most MCTS uses, random rollouts are performed until the end in order to estimate a so-called empirical mean. In our case, as we have already learned a value function for the states in the learning phase, we can directly plug this value as an initial estimate. Estimating the Q-value during selection also requires using a value for unexplored children nodes. Traditional MCTS uses an infinite value that forces in-breadth exploration of children. This is detrimental when the simulation budget is low wrt. the action space, as is in EOSP tasks. Here we copy the parent’s estimated value to the unexplored children (DBLP:journals/corr/abs-2104-04278).
@@ 5 Experiments
We use a set of real-life non-public problems consisting of 100 to 1500 candidate acquisitions to be scheduled over a single orbit. From those, there’s room for only about 50 of them to be selected for execution, yielding a largely over-subscribed problem for every orbit. As explained in Section 3, problems are given in the form of discrete graphs. The simulator uses this graph to compute and maintain the continuous-time graph. To provide intuition on the difficulty of the problem, Table 1 shows some statistics on a few test problems and their representation as graphs.
@@
# Acquisitions Discrete graph Continuous graph Ratio Discrete graph Continuous graph Ratio
@@
Table 1: Representative problem sizes
We compare our DRL approach, Wheatley, to two solutions currently being used for operating such satellites.
Greedy algorithm
It is the algorithm currently used for real-world operations. It greedily selects acquisitions to add to the schedule based on their utility, and inserts them in the plan if possible. Previously selected tasks may be slightly postponed, but never canceled.
RAMP
(RAMP): It is an implementation of a Dijkstra search algorithm in the discrete graph. Although based on an admissible algorithm (Dijkstra), RAMP is not guaranteed to find the absolute optimal schedule. This is due to the exclusion links between nodes of the discrete graph that represent different start dates of the same task. Nevertheless, RAMP constantly provides the best known schedules on real problems. Unfortunately, its complexity prevents using it for real-time operations. Therefore, it is used as a reference in these simulation results.
@@ 5.1 Unitary Score
First, we compare our approach to baselines on a relaxed problem: we try to maximize the number of acquisitions scheduled, irrespectively of their priority or utility. This measure of performance is insensitive to the large gaps in acquisitions utility discussed in Section 4.1. We run two experiments:
Single problem
First, we want to measure if our models and algorithms can possibly achieve competitive performance on a given problem. We train our learner on a single problem with a total of 106 acquisitions and let it overfit as much as needed, as long as it achieves great performance. As illustrated in Fig. 3, we observe that it is indeed able to outperform both the greedy algorithm and RAMP scores. This result shows that our architecture is able to implement very powerful policies. In the next set of experiments, we put it to the challenge of a realistic learning environment.
Figure 3: Unitary scores: single problem of 106 acquisitions.
Figure 4: Unitary scores: performance on 31 unseen problems (training on 128 different problems)
Generalization
To measure the ability to transfer knowledge from one task to another, we train on 128 problems of about 100 acquisitions and test on 31 unseen problems of similar size. The learning curve of Fig. 4 shows the evolution of the performance on the test set, as learning progresses. It peaks at around 600 training steps before slowly decreasing due to overfitting. We also measure the number of times where Wheatley’s performances are above, below or equal to the greedy algorithm (Fig. 5) and RAMP (Fig. 6) on the 31 test problems. This shows that our system is able to generalize to unseen problems, outperforming the currently deployed solution.
Figure 5: Unitary scores: Wheatley vs. Greedy
Figure 6: Unitary scores for Wheatley vs. Ramp
@@ 5.2 General Utility
In our second set of experiments, we take into account the utility of observations and aim at maximizing the cumulative utility of all the observations included in the final schedule, as in the full-fledged MDP framework presented in Section 4.1. As before, we perform two sets of experiments: one where the learner is free to overfit on a single problem to reach its best performance, and one aiming at measuring its ability to generalize.
Single Problem
Our test on a single problem with a total of 88 candidate acquisitions shows that our system is able to outperform the greedy algorithm and reach the score of RAMP (Figure 7). This proves the suitability of the architecture for the full MDP set up.
Figure 7: Utilities obtained when training on a single problem with 88 acquisitions
Figure 8: Utilities obtained averaging on 27 unseen problems after a training on 639 different problems
Generalization
We train on 639 problems of about 100 acquisitions and test on 27 unseen problems of similar size. The learning curves are displayed in Figure 8 and show that the learner is able to generalize. The plot showing the number of times where Wheatley is above, below or equal to the greedy solution are presented in Fig. 9 and same for RAMP are presented in Fig. 10. We see that Wheatley outperforms the deployed solution and approaches the best known performances, in a realistic set-up where problems are not known in advance. Fig. 11 shows examples of satellite trajectories produced by Greedy and Wheatley. It represents the roll angle (X-axis) at each acquisition date (Y-axis). Acquisition requests are depicted in gray during their VTW. Manoeuvers between acqusitions are the colored arrows describing the satellite depointing angle along time. We can see that Wheatley finds a smoother and more efficient sequence for the satellite.
Table 2 shows comprehensive results for the agent trained on problems of size 100, evaluated on different sizes of problems. The last line is an evaluation on instances with many conflicts, where RAMP performs worse than the greedy algorithm. Results show that Wheatley performs very well on not too large instances but is outperformed by the greedy approach on the largest instance. However, it is quite competitive in the case of highly conflictual instances, which is promising for future works.
Figure 9: Utilities for Wheatley vs. Greedy
Figure 10: Utilities for Wheatley vs. Ramp
@@
5.3 Inference Time MCTS
Results of policy search after learning is shown on Table 3, though on a different dataset [^4] The satellite target was updated during the work with support for a larger set of priority levels. than in Table 2. In this new dataset, the average number of selectable acquisitions is about 15, but the number of priority levels is higher, leading to larger operational scores when selecting high priorities. The displayed scores are the mean over the test set.
The budget is the number of trials done (i.e. the number of paths leading to tree expansion and backtrack of the estimates along the path) before selecting an action. MCTS uses the learned value network to get initial evaluation of new nodes. Two criterions are of interest: the mean absolute score on the test set, and the number of times where the MCTS search is better than RAMP reference. The results show that using MCTS at resolution times consistently improves found solutions over “vanilla” Wheatley.
It is worth noting that such budgets are small compared to the problem sizes, as the problems exhibit a large branching factor. For instance, in a problem of 100 acquisitions, a budget of 100 allows only tree exploration until a depth of 6, which is very far from the maximal depth of the tree (which could be up to 100 if all acquisitions are selectable). We are limited to such small budgets in order to perform search phase in sensible time: due to the update of discrete graphs and computations of continuous graphs at every node, this search phase duration is far from negligible, as shown in the results table.
score ($\times 10^{17}$) $\geq$ RAMP (%) time (min)
RAMP 8.62747 100
Wheatley 8.38172 28.7356
MCTS(100) 8.45111 37.9310 138
MCTS(1000) 8.56566 43.6782 1248
Table 3: MCTS(budget) results
@@ 6 Conclusion
We showed that DL-based approaches to the EOSP challenge some of the best known techniques. There are several perspectives we are currently exploring to extend this work. First, as stated before, we are trying to take advantage of the large discrepancy in acquisition utility by using lexicographic RL algorithms such as (lexicoRL). Scheduling tasks by decreasing priority would provide stronger guarantees to find the optimal schedule. To achieve this, schedules must be built in a non-chronological order, which is not the case in our current implementation. Currently, we choose the next acquisition to insert just after the last inserted one, using some foresight given by the GNN. This foresight is limited by the number of layers of the GNN. As we said, the discrete-time graph is tailored for state-space search and chronological insertion. Future work will consider developing an alternative continuous-time graph representation of the EOSP where observations can be added to the schedule in any order, using Simple Temporal Networks (stn). Such work will open promising avenues for using lexicographic preferences.